Find an equation of the line. Write the equation in the standard form. Through (6,8); parallel to 9x-y=6
Answer by mananth(16949)

9 x + -1 y = 6
Find the slope of this line
make y the subject
-1 y = -9 x + 6
Divide by -1
y = 9 x -6
Compare this equation with y=mx+b
slope m = 9
The slope of a line parallel to the above line will be the same
The slope of the required line will be 9
m= 9 ,point ( 6 , 8 )
Find b by plugging the values of m & the point in
y=mx+b
8 = 54.00 + b
b= -46
m= 9
Plug value of the slope and b in y = mx +b
The required equation is y = 9 x -46
y = 9 x -46
m.ananth@hotmail.ca